OpenStudy (anonymous):

sand is poured on a conical pile at a rate of 20 m^3/min. The height of the pile is always equal to the radius of the base of the pile. When the pile is 3m high, how fast is the height of the pile increasing?

imqwerty (imqwerty):

dv/dt=20 m^3/minute we r given that radius=height, so radius=3=heoght :) Volume of cone=1/3(πr^2h) now Volume=1/3π9h dv/dt=1/3(9)dh/dt dv/dt=3dh/dt dv/dt=3dh/dt we know that dv/dt=20 so 20=3dh/dt dh/dt=20/3

OpenStudy (anonymous):

how did you get the now volume?

OpenStudy (anonymous):

oh it's radius sorry :))

OpenStudy (anonymous):

*the

imqwerty (imqwerty):

it is given in the question that height of pike is ALWAYS equal to radius of pile :)

imqwerty (imqwerty):

*pile

OpenStudy (anonymous):

why was the π gone in dv/dt=1/3(9)dh/dt? :)

OpenStudy (anonymous):

@imqwerty

imqwerty (imqwerty):

yea pi shuld also come sry

OpenStudy (anonymous):

the answer should be 20/3π? :)

imqwerty (imqwerty):

yea

OpenStudy (anonymous):

oh it's not in the choices :(

imqwerty (imqwerty):

what r the choices?

OpenStudy (anonymous):

the choices are : 1.42, 0.71, 0.66 and 1.32

imqwerty (imqwerty):

something went wrng jst a min

OpenStudy (anonymous):

oh it's supposed to be 20/9π :)

OpenStudy (anonymous):

dv/dt = 1/3πr^(3)

imqwerty (imqwerty):

how?

OpenStudy (anonymous):



OpenStudy (anonymous):

dv/dt = πr^(2)dh/dt

imqwerty (imqwerty):

:O okay :)

OpenStudy (anonymous):

dv/dt = π(9) dh/dt then plug in 20 for dv/dt

OpenStudy (anonymous):

thank you guys for helping me :)

imqwerty (imqwerty):

no prblm :D